Write the balanced equation for the reaction between hydrochloric acid and magnesium. What mass of the magnesium product is formed when 133.77 g of hydrochloric acid reacts? Show work. Include Molar Mass Values and Dimensional Analysis.


1
Expert's answer
2021-03-02T02:00:29-0500

Mg + 2HCl --> MgCl2 + H2

Molar masses:

"M(HCl)=35.453+1.008=36.461g\/mol"

"M(MgCl_2)=24.305+2\times35.453=95.211g/mol"

Mass of MgCl2:

"m(MgCl_2)=133.77g(HCl)\times\frac{1mol(HCl)}{36.461g(HCl)}\times\frac{1mol(MgCl_2)}{2mol(HCl)}\times\frac{95.211g(MgCl_2)}{1mol(MgCl_2)}=174.66g"

Answer: 174.66 g